# Towards Human-AI Complementarity with Predictions Sets

### v1

@@ Appendix D Empirical average test accuracy on additional classification tasks
Figure 4: Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with conformal prediction (Naive and Aps) and our greedy algorithm (Greedy).
Appendix F Evaluation of the Mixture of Multinomial Logit Models (MNLs)
@@
Figure 5 shows that the average accuracy achieved by a simulated expert following the mixture of MNLs and by real human experts using the prediction sets constructed with all possible conformal predictors, each with a different $\alpha$ value, using the choice of calibration set by Straitouri et al. [19]. The results show that, while the mixture of MNLs tend to overestimates the average accuracy achieved by the predictions made by real experts, the average accuracy follows the same qualitative trend.
@@
Appendix G Conformal Prediction under different $\alpha$ values
In this section, we estimate the average accuracy achieved by a (simulated) expert using the prediction sets constructed with conformal prediction (Naive and Aps) on the ImageNet16H dataset under different $\alpha$ values. Figure 6 summarizes the results.
@@ Appendix G Conformal Prediction under different $\alpha$ values
Figure 6: Average accuracy achieved by a (simulated) expert using the prediction sets constructed with conformal prediction (Naive and Aps) on the ImageNet16H dataset under different $\alpha$ values. Each panel shows the average and standard error over $10$ runs.

### v2

@@ Appendix E Empirical average test coverage of prediction sets across classification tasks
Figure 4: Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with conformal prediction (Naive, Aps, Raps and Saps) and our greedy algorithm (Greedy).
Appendix G Evaluation of the Mixture of Multinomial Logit Models (MNLs)
@@
Figure 5 shows that the average accuracy achieved by a simulated expert following the mixture of MNLs and by real human experts using the prediction sets constructed with all possible conformal predictors, each with a different $\alpha$ value, using the choice of calibration set by Straitouri et al. [19]. The results show that, while the mixture of MNLs tends to overestimate the average accuracy achieved by the predictions made by real experts, the average accuracy follows the same qualitative trend.
@@
Appendix H Conformal Prediction under different $\alpha$ values
In this section, we estimate the average accuracy achieved by a (simulated) expert using the prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps) on the ImageNet16H dataset under different $\alpha$ values. Figure 6 summarizes the results.
@@ Appendix H Conformal Prediction under different $\alpha$ values
Figure 6: Average accuracy achieved by a (simulated) expert using the prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps) on the ImageNet16H dataset under different $\alpha$ values. Each panel shows the average and standard error over $10$ runs. We highlight with a red marker the highest average accuracy for the simulated humans under each conformal predictor.